The lifespan of the lightbulb in your bedroom follows an exponential distribution with a mean of ½ years. What is the expected number of lightbulbs to be burnt out in two years?
Answer #1

Mean = 1/2 years

So expected number of bulbs burnt in an year = 1/(1/2) = 2

Hence,

Expected number of lightbulbs to be burnt out in two years = 2*2 = 4